Find f(x) and find the equation of the line tangent to the graph off at x = 1. f(x) = (2+4x)(5 - 4x) f(x)=0 y=Find the x and y coordinates of all inflection points. f(x) = x3 + 18x? What is/are the inflection point(s)? Select the corre
Answer #1

f(x) = (2+4+) (5-4x) f(x) = 4 C5_407+ (-4) (2+ 4 s) 20-16- 8 - 162 12-322 f( 12-32(1) -20 f(12= (2+4) (5-4) = 6*1 = 6 (rigg